how do i solve (r+6)/(5)=(r+8)/(7)

Answer by stanbon(75887)

solve (r+6)/(5)=(r+8)/(7)
-------
Multiply both sides by 35 to get:
7(r+6) = 5(r+8)
---
7r+42 = 5r+40
Subtract 5r from both sides.
Subtract 42 from both sides.
2r = -2
r = -1
===================
Cheers,
Stan H.